A random sample of 300 random people from a geographic area indicated that 50 of 

them had received vaccine. Estimate the true proportion of people who had been vaccinated with 95% confidence.


1
Expert's answer
2021-05-07T14:25:48-0400

Solution:

"\hat p=\frac{50}{300}=\frac16"

"n=300"

"\hat q=1-\frac16=\frac56"

For 95% confidence interval, "z=1.96"

Now, 95% confidence interval"=(\hat p\pm z\sqrt{\frac{\hat p \hat q}{n}})"

"=(\frac16\pm 1.96\sqrt{\frac{\frac16. \frac56}{300}})
\\=(0.1244,0.2088)"